# Heterogeneous Multi-Agent Proximal Policy Optimization for Power Distribution System Restoration

### v4

@@ III-E Training Workflow
Algorithm 1 summarizes the unified HAPPO training procedure used for both IEEE 123-bus and IEEE 8500-node feeders. Each training iteration consists of several steps: agents first interact with OpenDSS for $T$ steps to generate rollouts; the centralized critic then computes temporal-difference errors and $\lambda$-GAE advantages; each agent subsequently updates its actor parameters through sequential clipped PPO-style optimization; the critic is regressed against bootstrapped returns to refine value estimates; and finally, the updated actor and critic networks are deployed for the next episode. This unified workflow enables the method to scale seamlessly between medium- and large-scale feeders, ensuring stable, reproducible learning dynamics.
1
@@
2
3Initialize: agents, critic $V_{\phi}$, rollout buffer $\mathcal{B}$, and logger;
4 Load hyperparameters: rollout length $T$, learning rates, clipping threshold $\epsilon$, discount $\gamma$, GAE parameter $\lambda$, etc.;
5
6for each iteration $u=1$ to $U_{\max}$ do
7  Reset environment and obtain $(o_{k,0},s_{0})$;
8
@@ 7  Reset environment and obtain $(o_{k,0},s_{0})$;
10     foreach agent $k$ do
11        Sample action $a_{k,t}\sim\pi_{\theta_{k}}(a|o_{k,t})$;
12
@@ 15     Store transition in buffer $\mathcal{B}$;
17
18 Compute TD errors and GAE advantages: $\delta_{t}=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),\quad\hat{A}_{t}=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\delta_{t+\ell}$
19 foreach agent $k$ (sequential update) do
20     Freeze $\pi_{\theta_{j}}$ for all $j\neq k$;
@@
22        Sample minibatches $(o_{k},a_{k},\hat{A},\log\pi_{\text{old}})$;
23        Compute ratio $r$ and clipped surrogate objective;
24        Update actor parameters $\theta_{k}\leftarrow\theta_{k}-\alpha_{\pi}\nabla L_{\text{actor}}$;
25
26
27
28 Compute critic loss: $L_{\text{critic}}=\mathbb{E}\!\left[(V_{\phi}(s_{t})-(\hat{A}_{t}+V_{\phi}(s_{t})))^{2}\right]$ Update critic: $\phi\leftarrow\phi-\alpha_{V}\nabla L_{\text{critic}}$;
29
30 Log training metrics and save checkpoints;
31
@@ 30 Log training metrics and save checkpoints;
33return trained $\{\pi_{\theta_{k}}\}_{k=1}^{N}$ and critic $V_{\phi}$;
@@ IV-A Simulation Environment
The IEEE 123-bus feeder (Fig. 2) contains 123 buses, a 3025 kW total load, five DER units with 2400 kW combined capacity, and 26 controllable switches. For multi-agent training, it is divided into five microgrids, each controlled by one agent responsible for local switching. The feeder exhibits realistic unbalanced loading and DER interactions.
@@
The IEEE 8500-node feeder contains roughly 8500 nodes, a 25 MW total load, 100 controllable switches, and DERs providing up to 2400 kW. It is partitioned into ten microgrids, each managed by a dedicated agent. The large state dimensionality and deep radial topology make it a challenging testbed for evaluating scalable multi-agent coordination with HAPPO.
@@ IV-B Training Configuration
All experiments follow the unified HARL training workflow in Section III (Algorithm 1). Each episode runs for $T$ steps, with decentralized recurrent actors producing discrete switching actions and a centralized critic $V_{\phi}(s_{t})$ computing values and $\lambda$-GAE advantages. Policy learning uses PPO-style updates with clipping threshold $\epsilon$, entropy coefficient $\beta_{\text{ent}}$, and sequential per-agent optimization. Rewards include incremental restored power, normalized loss penalties, and continuous constraint-violation terms. A fixed generation cap of $P_{\text{gen}}=2400\,\text{kW}$ is applied to both feeders. Hyperparameters ($\gamma$, $\lambda$, learning rates, batch sizes, etc.) remain identical across systems unless scale adjustments are required.
IV-C Hardware and Implementation Details
Training is performed on a workstation with an NVIDIA RTX 6000 Ada GPU (48 GB), AMD Threadripper PRO CPU, and 256 GB RAM. OpenDSS runs on the CPU, while PyTorch handles actor–critic updates on the GPU. The HARL framework supports decentralized actors, a centralized critic, sequential PPO updates, and synchronized rollout collection.
IV-D Evaluation Metrics
Performance is evaluated using cumulative reward (weighted restoration benefit), restored active power (relative to the 2400 kW cap), convergence stability across seeds, inference latency per switching action, and total training time. This unified protocol enables consistent comparison across feeders and baseline RL/MARL methods.
@@ V Numerical Results
This section evaluates the proposed HAPPO-based restoration framework on the IEEE 123-bus and IEEE 8500-node systems using the unified HARL training pipeline described in Sections IIIIV. Performance is reported in terms of (i) cumulative reward, which reflects the weighted restoration benefit accumulated during each episode, and (ii) restored active power, which directly measures physical restoration effectiveness relative to the system-wide generation cap $P_{\text{gen}}=2400\,\text{kW}$. All results are averaged over multiple independent random seeds.
V-A Training Performance on IEEE 123-Bus and IEEE 8500-Node Systems
Figures 3 and 4 show the evolution of cumulative reward and restored active power during training for the two feeders. Each curve corresponds to an independent seed initialized with randomized outage locations, critical-load placements, and DER conditions.
On both feeders, agents select among three discrete switching actions (open, close, or no-op) and receive a shared reward that incorporates incremental restored power, normalized loss penalties, and differentiable constraint-violation terms. Unlike masking-based MARL methods, the proposed penalty-driven reward ensures uninterrupted learning near feasibility boundaries and supports smooth gradient propagation.
In the IEEE 123-bus feeder 5 agents coordinate 26 switches. Across five independent training runs, the cumulative reward increases steadily and converges to $(3.6\pm 2.9)\times 10^{6},$ while the restored power stabilizes at $2294\pm 52~\text{kW}\quad(95.6\%\pm 2.2\%\text{ of }P_{\text{gen}}).$ As shown in Fig. 3, performance across seeds is tightly clustered, indicating reliable convergence and stable sequential actor updates.
In the IEEE 8500-node feeder, 10 agents each govern ten randomly assigned switches across 1,177 loads. Despite the dramatically larger state–action space and deeper network topology, HAPPO maintains monotonic improvement. The cumulative reward converges to $(2.9\pm 3.3)\times 10^{6},$ and restored power reaches $2308\pm 55~\text{kW}\quad(\!96.2\%\pm 2.3\%\text{ of }P_{\text{gen}}),$ as shown in Fig. 4. These results highlight the scalability of sequential trust-region updates, centralized advantage estimation, and penalty-based reward shaping.
Figure 3: Training performance of HAPPO on the IEEE 123-bus feeder across four random seeds. Left: cumulative reward. Right: restored active power.
Figure 4: Training performance of HAPPO on the IEEE 8500-node feeder across five random seeds. Left: cumulative reward. Right: restored active power.
TABLE I: Performance comparison of HAPPO with state-of-the-art RL/MARL methods on IEEE 123-bus and IEEE 8500-node feeders (generation cap $P_{\text{gen}}=2400\,\text{kW}$)
@@ V-A Training Performance on IEEE 123-Bus and IEEE 8500-Node Systems
Cum. Reward Restored (%) Train Time (min) Inference (ms) Cum. Reward Restored (%) Train Time (min) Inference (ms)
DQN [13] 28.38 $\pm$ 0.14 73.90 $\pm$ 0.46 34.7 4.2 65.58 $\pm$ 0.25 69.50 $\pm$ 0.28 249.3 6.8
PPO [27] 29.01 $\pm$ 0.15 74.85 $\pm$ 0.47 44.8 7.1 67.94 $\pm$ 0.35 71.05 $\pm$ 0.27 255.7 9.8
MAES [17] 29.78 $\pm$ 0.19 76.14 $\pm$ 0.40 91.8 25.4 69.82 $\pm$ 0.32 72.81 $\pm$ 0.23 307.0 38.7
MAGDPG [18] 30.08 $\pm$ 0.17 77.00 $\pm$ 0.42 118.2 29.5 70.56 $\pm$ 0.24 73.44 $\pm$ 0.32 321.5 45.6
MADQN [24] 30.58 $\pm$ 0.22 77.80 $\pm$ 0.59 98.8 12.6 71.37 $\pm$ 0.34 74.02 $\pm$ 0.39 279.2 19.0
Mean-Field RL [19] 31.56 $\pm$ 0.25 80.26 $\pm$ 0.52 108.3 27.3 73.46 $\pm$ 0.33 77.40 $\pm$ 0.38 303.9 42.6
QMIX [20] 33.38 $\pm$ 0.25 83.25 $\pm$ 0.45 96.4 15.8 76.62 $\pm$ 0.34 80.10 $\pm$ 0.30 291.3 23.5
[…]
@@ V-B Multi-Seed Convergence and Reproducibility
On the IEEE 123-bus feeder, restored power converges within $2230\text{--}2367~\text{kW},$ corresponding to 93–99% of the generation cap. On the IEEE 8500-node feeder, restored power stabilizes near $2224\text{--}2347~\text{kW},$ representing 93–98% of the available capacity.
@@ V-C Comparative Evaluation with Baseline Methods
HAPPO achieves the highest restored power and cumulative reward on both feeders. Its sequential per-agent trust-region updates and centralized advantage estimation enable stable policy optimization even in the presence of strong physical coupling and high-dimensional state spaces, outperforming value-based and value-decomposition methods that struggle under these conditions.
@@ V-D Discussion and Interpretation
The results demonstrate four key findings. (1) HAPPO exhibits strong scalability, extending from a five-agent IEEE 123 system to a ten-agent IEEE 8500 system with minimal degradation in convergence smoothness. (2) The differentiable penalty-based reward enables continuous guidance near operational limits, avoiding the brittleness typically observed in masking-based approaches and enhancing constraint awareness. (3) Multi-seed experiments show consistently low variance, confirming the reproducibility of sequential PPO-style updates. (4) The OpenDSS-based simulation environment ensures physical fidelity by enforcing operational limits and producing policies that correspond to feasible feeder configurations in realistic distribution systems.
These findings confirm that the HARL-based HAPPO framework is well suited for large-scale, constraint-driven distribution system restoration.
@@ VI Conclusion
This paper presented a HAPPO approach for distribution system restoration within a HARL framework using an OpenDSS-based multi-agent environment. Sequential trust-region updates combined with centralized advantage estimation enable stable, sample-efficient coordination among structurally heterogeneous microgrid agents, avoiding the instability issues common in conventional MARL. A continuous penalty-based reward replaces masking or early termination, yielding smooth gradients and naturally constraint-aware switching behavior.
Experiments on the IEEE 123-bus and IEEE 8500-node feeders show that HAPPO consistently outperforms PPO, QMIX, Mean-Field RL, and other baselines. Under the 2400 kW generation cap, it restores roughly 92% and 96% of available load on the two systems, respectively, with strong multi-seed reproducibility and low inference latency. The OpenDSS integration ensures physical validity and supports direct transition to operational settings.
Overall, the results demonstrate that heterogeneous-agent policy optimization provides a scalable, stable, and realistic solution for resilient distribution system restoration. Future work will incorporate model-based elements, communication constraints, partial observability, and dynamic DER participation to further strengthen real-time applicability.

### v5

@@ III-E Training Workflow
Algorithm 1 summarizes the unified HAPPO training procedure for both IEEE 123-bus and 8500-node feeders. Each iteration involves: agents interact with OpenDSS for $T$ steps to generate rollouts; the centralized critic computes temporal-difference errors and $\lambda$-GAE advantages; agents sequentially update actor parameters via clipped PPO-style optimization; the critic regresses against bootstrapped returns to refine value estimates; and updated networks deploy for the next episode. This workflow enables seamless scaling between medium- and large-scale feeders, ensuring stable, reproducible learning dynamics.
@@
1
21exInitialize: agents, critic $V_{\phi}$, rollout buffer $\mathcal{B}$, and logger;
3 Load hyperparameters: rollout length $T$, learning rates, clipping threshold $\epsilon$, discount $\gamma$, GAE parameter $\lambda$, etc., and checkpoint interval $S$ (save period);
4
51exfor each iteration $u=1$ to $U_{\max}$ do
6  Reset environment and obtain $(o_{i,0},s_{0})$;
@@
8     foreach agent $i$ do
9        Sample action $a_{i,t}\sim\pi_{\theta_{i}}(a|o_{i,t})$;
10
@@ 13     Store transition in buffer $\mathcal{B}$;
14
15 Compute TD errors and GAE advantages for each agent: $\delta_{t}^{i}=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),\quad\hat{A}_{t}^{i}=\sum_{\ell=0}^{T-t-1}(\gamma\lambda)^{\ell}\delta_{t+\ell}^{i}$;
@@
17 foreach agent $i$ (sequential update) do
18     Freeze $\pi_{\theta_{j}}$ for all $j\neq i$;
@@
20        Sample minibatches $(o_{i},a_{i},\hat{A},\log\pi_{\text{old}})$;
21        Compute probability ratio $\rho$ and clipped surrogate objective;
22        Update actor parameters $\theta_{i}\leftarrow\theta_{i}-\alpha_{\pi}\nabla L_{\text{actor}}$;
23
24
25 Compute critic loss: $L_{\text{critic}}=\mathbb{E}\!\left[(V_{\phi}(s_{t})-\hat{R}_{t})^{2}\right]$,
26  where $\hat{R}_{t}=\sum_{\ell=0}^{T-t-1}\gamma^{\ell}r_{t+\ell}+\gamma^{T-t}V_{\phi}(s_{T})$;;
27  Update critic: $\phi\leftarrow\phi-\alpha_{V}\nabla L_{\text{critic}}$;
28
29 Log training metrics at iteration $u$ (iter);
30  if $u\bmod S=0$ then
31     Save model checkpoint (ckpt);
@@
33
34return trained $\{\pi_{\theta_{i}}\}_{i=1}^{N}$ and critic $V_{\phi}$;
@@ IV-A Simulation Environment
The IEEE 123-bus feeder (Fig. 2) has 123 buses, a 3025 kW load, five DERs totaling 2400 kW, and 26 controllable switches, partitioned into five agent-controlled microgrids. The IEEE 8500-node feeder contains about 8500 nodes, a 25 MW load, 100 switches, and 2400 kW of DER capacity, divided into ten microgrids. Its scale and deep radial structure make it a demanding testbed for HAPPO scalability.
@@ IV-B Training Configuration
All experiments follow the unified HARL workflow in Section III (Algorithm 1). Each $T$-step episode uses decentralized recurrent actors for discrete switching actions and a centralized critic $V_{\phi}(s_{t})$ for values and $\lambda$-GAE advantages. Policy learning employs PPO-style updates with clipping threshold $\epsilon$, entropy coefficient $\beta_{\text{ent}}$, and sequential per-agent optimization. Rewards combine incremental restored power, normalized loss penalties, and continuous constraint violations. Both feeders use a fixed generation cap $P_{\text{gen}}=2400\,\text{kW}$. Hyperparameters ($\gamma$, $\lambda$, learning rates, batch sizes) remain identical unless scale adjustments are needed.
IV-C Hardware and Evaluation Metrics
Training is performed on a workstation with an NVIDIA RTX 6000 Ada GPU (48 GB), AMD Threadripper PRO CPU, and 256 GB RAM, with OpenDSS on CPU and PyTorch on GPU. Performance is evaluated using restored active power, convergence stability across seeds, inference latency, and training time, enabling consistent comparison across feeders and baseline RL/MARL methods.
@@ V Numerical Results
V-A Performance on IEEE 123- and 8500-Bus Systems
Figures 3 and 4 show the evolution of restored active power during training for the two feeders. Each curve corresponds to an independent seed initialized with randomized outage locations, critical-load placements, and DER conditions.
In the IEEE 123-bus feeder, 5 agents coordinate 26 switches. Across five independent random seeds, the restored power stabilizes at $2294\pm 53~\text{kW}(95.6\%\pm 2.2\%\text{ of }P_{\text{gen}}).$ As shown in Fig. 3, performance across seeds is tightly clustered, indicating reliable convergence and stable sequential actor updates.
In the IEEE 8500-node feeder, 10 agents each govern ten randomly assigned switches across 1,177 loads. Despite the dramatically larger state–action space and deeper network topology, HAPPO maintains monotonic improvement. The restored power reaches $2309\pm 55~\text{kW}\quad(\!96.2\%\pm 2.3\%\text{ of }P_{\text{gen}}),$ as shown in Fig. 4. These results highlight the scalability of sequential trust-region updates, centralized advantage estimation, and penalty-based reward shaping.
Figure 3: Training performance of HAPPO on the IEEE 123-bus feeder across five random seeds showing restored active power evolution.
Figure 4: Training performance of HAPPO on the IEEE 8500-bus system across three random seeds showing restored active power evolution.
TABLE I: Performance comparison of HAPPO with state-of-the-art RL/MARL methods on IEEE 123-bus and IEEE 8500-node feeders
@@ V-A Performance on IEEE 123- and 8500-Bus Systems
Restored (%) Train Time (min) Inference (ms) Restored (%) Train Time (min) Inference (ms)
DQN [13] 73.90 $\pm$ 0.46 34.7 4.2 69.50 $\pm$ 0.28 249.3 6.8
PPO [27] 74.85 $\pm$ 0.47 44.8 7.1 71.05 $\pm$ 0.27 255.7 9.8
MAES [17] 76.14 $\pm$ 0.40 91.8 25.4 72.81 $\pm$ 0.23 307.0 38.7
MAGDPG [18] 77.00 $\pm$ 0.42 118.2 29.5 73.44 $\pm$ 0.32 321.5 45.6
MADQN [24] 77.80 $\pm$ 0.59 98.8 12.6 74.02 $\pm$ 0.39 279.2 19.0
Mean-Field RL [19] 80.26 $\pm$ 0.52 108.3 27.3 77.40 $\pm$ 0.38 303.9 42.6
QMIX [20] 83.25 $\pm$ 0.45 96.4 15.8 80.10 $\pm$ 0.30 291.3 23.5
HAPPO (ours) 95.6 $\pm$ 2.2 110.5 22.1 96.2 $\pm$ 2.3 391.2 33.4
@@ V-B Multi-Seed Convergence and Reproducibility
On the IEEE 123-bus feeder, restored power converges within $2241\text{--}2347~\text{kW},$ corresponding to 93–98% of the generation cap. On the IEEE 8500-node feeder, restored power stabilizes near $2254\text{--}2364~\text{kW},$ representing 94–99% of the available capacity.
@@ V-C Comparative Evaluation with Baseline Methods
HAPPO achieves the highest restored power on both feeders. Its sequential per-agent trust-region updates and centralized advantage estimation enable stable policy optimization even in the presence of strong physical coupling and high-dimensional state spaces, outperforming value-based and value-decomposition methods that struggle under these conditions.
@@ V-D Discussion and Interpretation
The results demonstrate four key findings. (1) HAPPO exhibits strong scalability, extending from a five-agent IEEE 123 system to a ten-agent IEEE 8500 system with minimal degradation in convergence smoothness. (2) The differentiable penalty-based reward enables continuous guidance near operational limits, avoiding the brittleness typically observed in masking-based approaches and enhancing constraint awareness. (3) Multi-seed experiments show consistently low variance, confirming the reproducibility of sequential PPO-style updates. (4) The OpenDSS-based simulation environment ensures physical fidelity by enforcing operational limits and producing policies that correspond to feasible feeder configurations in realistic distribution systems. These findings confirm that the HARL-based HAPPO framework is well suited for large-scale, constraint-driven distribution system restoration.
Although a full ablation study is beyond this work’s scope, benchmark methods lacking centralized critics or heterogeneous-agent modeling (e.g., independent PPO, parameter-sharing MAPPO) show reduced stability in strongly coupled feeders. The consistent multi-seed convergence in Figures 34 suggest centralized advantage estimation and microgrid-level heterogeneity are key contributors; systematic ablation remains for future work.
Training uses full AC power-flow simulation on a single GPU, as offline training is not time-critical. At deployment, policy inference runs in milliseconds (under 35 ms per step on the IEEE 8500-node feeder), meeting practical restoration requirements.
@@ VI Conclusion
This paper presented a HAPPO-based HARL framework for distribution system restoration using OpenDSS. Sequential trust-region updates with centralized advantage estimation enable stable coordination among heterogeneous microgrid agents, while continuous penalty-based rewards provide constraint-aware switching behavior.
Experiments on the IEEE 123-bus and 8500-node feeders show that HAPPO outperforms DQN, PPO, MAES, MAGDPG, MADQN, Mean-Field RL, and QMIX, restoring over 95% of available load under a 2400 kW cap with strong multi-seed stability and under 35 ms inference latency. Future work will explore model-based extensions, communication limits, partial observability, and dynamic DER participation.v